OpenStudy (1dfanatic):

Find the slope of a line parallel to the graph of -3x - 5y = 6

OpenStudy (jdoe0001):

if you solve that for "y", what does it look like?

OpenStudy (fibonaccichick666):

can you put your eq in slope-intercept form please?

OpenStudy (1dfanatic):

5y=-3x+6?

OpenStudy (fibonaccichick666):

so how do you move that 5 to the other side?

OpenStudy (1dfanatic):

subtract it?

OpenStudy (fibonaccichick666):

you also dropped your negative signs

OpenStudy (fibonaccichick666):

so let's take a step back

OpenStudy (fibonaccichick666):

how do you get -3x to the other side?

OpenStudy (1dfanatic):

add it?

OpenStudy (fibonaccichick666):

ok and what does that give you?

OpenStudy (1dfanatic):

Like the the equation or the answer?

OpenStudy (fibonaccichick666):

when you add 3x to both sides, what is the result?

OpenStudy (1dfanatic):

-3x-8y=9?

OpenStudy (fibonaccichick666):

uhm... where did that come from?

OpenStudy (1dfanatic):

idk im confused tbh..

OpenStudy (fibonaccichick666):

ok, so let's try it this way. you have \[-3x - 5y = 6\] when we add 3x to both sides it looks like this before simplifying \[3x-3x - 5y = 6+3x\]

OpenStudy (fibonaccichick666):

what is 3x-3x?

OpenStudy (1dfanatic):

0

OpenStudy (fibonaccichick666):

so what does the eq above simplify to?

OpenStudy (fibonaccichick666):

what's the next step?

OpenStudy (1dfanatic):

divide by 6?

OpenStudy (fibonaccichick666):

I just want the simplification

OpenStudy (fibonaccichick666):

what do you write after 3x−3x−5y=6+3x